OpenStudy (abdela25):

write the following equation in slope-intercept form... 6x+9y=27 medals will be awarded to first best answer

OpenStudy (anonymous):

Slope-intercept form is y = mx + b So we just basically get y = whatever it ends up being 6x+9y=27 Move over the x, by subtracting 6x from both sides. 9y = 27 - 6x Now let's divide it all by 9 and re-arrange y = -2/3 x + 3

OpenStudy (abdela25):

seems pretty good but im still confused medal was awarded further explaining will result in me being a fan @G-unit

OpenStudy (abdela25):

step by step i think would clarify things

OpenStudy (anonymous):

Okay, so you're given an equation, that has x's, y's, and numbers. The point slope form is y = mx + b Y stands for Y, or 1Y. m stands for the slope, X stands for X, and b stands for the intercept. None of that matters though, in this question. We just have to make sure we end up having Y as Y, and nothing else. That means we can't leave it as 9Y. So let's start again, and this time divide everything by 9 first. 9Y divided by 9 is Y 6X divided by 9 is 6x/9, which we can simplify to 2X/3 by dividing both top and bottom by 3, right? Then 27 divided by 9 is 3. Now let's re-visualize it 2X/3 + Y = 3 Let's get the Y on the left side by itself now. We'll just subtract the whole X part to the other side. -2X/3 on both sides. We get Y = -(2X/3) + 3 We can rewrite the X bit as... -2/3 X So now we have Y = -2/3 X + 3

OpenStudy (anonymous):

By the way, it would be (-2/3)x and not (-2)/(3x) So it would look like \[\frac{ 2 }{ 3 } x\]

OpenStudy (abdela25):

thank you very much!!! greatly appreciated nice @G-unit